Calculus II, Spring 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 12

1. Section 14.7: Problems 3, 5, 8 (can you figure out how one chooses the change of variables u, v in this problem?), 11, 13.

2. Chap 14: 28, 33, 37, 39, 41 (need not evaluate).